For the balanced equation shown below, how many grams of C2H3O2Cl will react with 14.5 grams of HCl? 2C2H3O2Cl(g) + 3O2 (g) -->4CO2 (g) + 2H2O(g) + 2HCl (g)
Answer #1

Molar mass of HCl,

MM = 1*MM(H) + 1*MM(Cl)

= 1*1.008 + 1*35.45

= 36.458 g/mol

mass of HCl = 14.5 g

mol of HCl = (mass)/(molar mass)

= 14.5/36.46

= 0.3977 mol

According to balanced equation

mol of C2H3O2Cl = moles of HCl

= 0.3977 mol

Molar mass of C2H3O2Cl,

MM = 2*MM(C) + 3*MM(H) + 2*MM(O) + 1*MM(Cl)

= 2*12.01 + 3*1.008 + 2*16.0 + 1*35.45

= 94.494 g/mol

mass of C2H3O2Cl = number of mol * molar mass

= 0.3977*94.49

= 37.58 g

Answer: 37.6 g